OpenStudy (anonymous):

\[\int\limits_{?}^{?}(1 + 1/x^2 +1/x^3) dx\] help pls...

OpenStudy (anonymous):

\[\int\limits 1 + x ^{-2} ^{-3}\]

OpenStudy (anonymous):

\[\int\limits 1 +x ^{-2} +x ^{-3}\]

OpenStudy (anonymous):

\[x +x ^{-2+1}/(-2+1) +x ^{-3+1}/(-3+1)\]

OpenStudy (anonymous):

x -1/x-1/2x^2

OpenStudy (anonymous):

intergral by parts, \[\int\limits_{}^{} 1 dx = x\] \[\int\limits_{}^{} x^{-2} = x^{-2+1}/(-2+1) =- x^{-1}\] \[\int\limits_{}^{} x^{-3} = x^{-3+1}/(-3+1) =- x^{-2}/2\]

OpenStudy (anonymous):

tnx a lot uzma ...:)

OpenStudy (anonymous):

tnx amogh!!